# Energy Weighted Learning Progress Guided Interleaved Multi-Task Learning

### v1

@@ V-C Overall Learning Performance
Figure 3: Task-specific model performance is shown. It can be seen that LP-based method surpasses the other baselines on both push and stack tasks, with a higher significance in the latter. However, for the hit task, the performance is not improved considering the other two baselines.
V-D Network Complexity
Figure 4: Model performances with different neural network complexity levels. Even though at high parameter regimes, all models perform similarly, the LP-based learning method (INTER-LP) becomes superior at low or mid-parameter regimes, where suitable task switching for appropriate skill transfer among tasks becomes crucial.
To analyze the impact of network complexity on performance, we evaluate baselines (SINGLE and INTER-RAND) and our proposed model (INTER-LP) as follows: Each model is trained with three levels of network complexity: Low ($\sim 800$ parameters), Medium ($\sim 2000$ parameters), and High ($\sim 5200$ parameters). The results, as shown in Figure 4, illustrate how performance evolves across different resource levels.
Low Complexity. At the lowest network complexity, INTER-LP achieves a significantly lower MAE compared to the two baselines, indicating its ability to efficiently utilize limited resources. This result highlights the effectiveness of our learning progress-based interleaving strategy in scenarios where computational resources are constrained. In contrast, SINGLE and INTER-RAND exhibit higher prediction errors, suggesting that they are less capable of optimizing task performance when resources are limited.
Medium Complexity. As the network size increases, the performance gap between INTER-LP and the two baselines begins to narrow. While INTER-LP still outperforms both SINGLE and INTER-RAND, the relative improvements are less pronounced compared to the low-complexity scenario. This indicates that the baselines start benefiting from the additional parameters, but INTER-LP continues to leverage its task-interleaving strategy to maintain superior performance.
High Complexity. At the highest level of network complexity, the difference between INTER-LP and the baselines is further reduced, with all three models achieving relatively similar levels of performance. This result suggests that as the network’s capacity grows, both SINGLE and INTER-RAND can partially compensate for their limitations in interleaving strategy by relying on the increased capacity for task representation. However, INTER-LP remains competitive, showcasing better performance across all network complexities.
V-E Interleaved vs. Blocked Learning
@@ V-D Network Complexity
Figure 5: Interleaved learning versus blocked learning in all possible orders (titles showing the training order of the tasks). No matter the order in which the tasks are trained, the interleaved learning method always surpasses the blocked learning method, which shows that interleaved learning prevents catastrophic forgetting of learned information by switching the context in periods, thus being able to recall faster.
We conducted a detailed comparison of interleaved learning and blocked learning in a multi-task learning network across the three tasks (Push, Hit, and Stack). The interleaved learning mode, implemented in our model INTER-LP, alternates between tasks during training, while the blocked learning mode, implemented in the BLOCKED model, learns each task sequentially in its entirety before moving to the next.
Given the three tasks, there are six possible training orders (e.g., Push $\rightarrow$ Hit $\rightarrow$ Stack or Push $\rightarrow$ Stack $\rightarrow$ Hit, etc.). To see whether any order can perform similar or bettered that interleaved learning, we trained the BLOCKED model across all six task orders, while the INTER-LP model remained unchanged across these configurations, as it does not depend on the task order. Figure 5 shows the performance of the INTER-LP model and the BLOCKED model for each of the six task orders. It can be seen that the proposed model, INTER-LP consistently outperforms the BLOCKED model across all task orders. This consistent performance demonstrates the robustness of interleaved learning, as it mitigates the catastrophic forgetting observed in the Blocked model. Specifically, in blocked learning, the model exhibits a marked decrease in performance on earlier tasks as it progresses through subsequent tasks. For instance, when training starts with a Push task in the BLOCKED model, its performance on a Push task significantly degrades by the time Hit and Stack are learned. Additionally, the variability in performance across different task orders for the BLOCKED model highlights its sensitivity to the order of training. This variability further emphasizes the limitations of blocked learning in scenarios where task interdependence or order effects play a critical role.
V-F EWLP Based Task Selection Performance
Since we hold that the human brain also considers computational cost, in this section, we explore discounting the learning progress of the tasks by the neural cost they incur for task switching. We look at different discount levels and compare the learning performances. Figure 6 presents the overall learning performance of our proposed method (INTER-LP), its energy-weighted variants (INTER-LPE-K1 and INTER-LPE-K1.2), and the baseline models (SINGLE and INTER-RAND), in terms of prediction performance. As expected INTER-LP shows the lowest mean absolute error, with a marked performance margin during the early and mid stages. The energy-weighted variants, INTER-LPE-K1 and INTER-LPE-K1.2, also outperform both SINGLE and INTER-RAND baselines, indicating that incorporating energy awareness does not nullify the gains of our architecture. In fact, INTER-LPE-K1.2 shows performance nearly on par with INTER-LP, suggesting that slightly increasing the weight of learning progress relative to energy cost (via the sensitivity constant $k$) can further enhance training efficiency. The SINGLE model exhibits the weakest performance overall, reaffirming that task isolation limits skill learning efficiency. Meanwhile, INTER-RAND shows modest improvements over SINGLE, highlighting that even random task interleaving can offer some benefit, yet remains clearly inferior to structured LP-based scheduling. Taken together, these results confirm that both learning progress and energy-aware arbitration strategies provide meaningful gains in training efficiency and final task performance, with INTER-LP and INTER-LPE variants converging faster and reaching lower prediction errors than all baselines.
@@ V-F EWLP Based Task Selection Performance
Figure 6: Comparison of baselines and LP-based model to the EWLP-based models where the sensitivity constant is set to $k=1$ and $k=1.2$, respectively.
@@
To further evaluate the effect of EWLP-based task scheduling, we investigate the performance of the INTER-LPE model under varying values of the energy sensitivity coefficient $k$, and compare it to INTER-LP and SINGLE baselines. As shown in Figure 7a, which reports the MAE exactly at epoch 1000, INTER-LP achieves the lowest prediction error overall, confirming the strong effectiveness of LP-based task arbitration. Then, INTER-LPE-K1.2 closely follows, demonstrating that modest energy modulation can preserve much of the learning performance. As $k$ decreases, the prediction performance gradually degrades, indicating that overly prioritizing energy conservation (i.e., smaller $k$) can hinder learning effectiveness. On the other hand, Figure 7b shows total energy consumption measured in terms of cumulative neuron activations. Here, we observe the opposite trend: energy usage decreases steadily with smaller $k$, with INTER-LPE-K0.4 consuming the least energy overall, almost the same with the SINGLE baseline. While INTER-LP achieves the best predictive performance, it is also the most energy-intensive. Notably, INTER-LPE-K1.2 and K1 provide a favorable trade-off—offering significant reductions in energy consumption with only a small sacrifice in prediction accuracy. This tunable behavior highlights the flexibility of the INTER-LPE framework, enabling agents to balance performance and energy efficiency according to task demands and environmental constraints.
(a) Training performances of the models.
(b) Energy consumption of models.
@@
Figure 7: Effect of varying the energy sensitivity coefficient $k$ in the INTER-LPE model on (a) prediction performance and (b) total energy consumption. Higher values of $k$ prioritize learning progress over energy efficiency, leading to lower prediction error but increased energy usage, while lower $k$ values reduce energy consumption at the cost of predictive accuracy.
V-G Ablation Experiments
V-G1 Necessity of Attention and the Flag Bit
@@
Figure 8: Contributions of the shared attention and flag bit from the proposed multi-task learning architecture to the model performances. Neither attention nor flag alone can surpass the convergence speed of our proposed model, which combines them together.
@@
To evaluate the contributions of the attention layer and the flag bit in our proposed model INTER-LP, we conducted an ablation study by training four variations of the model under different configurations. Each model was trained using 10 different random seeds to ensure the robustness of the results. Below, we describe the four settings and provide a detailed analysis of their performance:
Ablation: none. This is our proposed model, which includes a shared attention layer taking the matrix $Z$ from (6), which is formed using the task representations along with their corresponding flag bits. The flag bits serve to differentiate tasks, while the attention layer dynamically learns task interactions.
@@
Ablation: flag. This variation retains the shared attention layer but removes the flag bits. Task representations are combined and passed to the shared attention layer without explicit differentiation using flags.
@@
Ablation: attention. In this model, the flag bits are concatenated to the task representations, the same as the original matrix $Z$ (6), but the shared attention layer is removed from the network.
@@
Ablation: attention + flag. This configuration removes both the shared attention layer and the flag bits. Only the row-wise concatenated task representations are passed to the task decoder of the current training task.
The results of the experiments that are shown in Figure 8 demonstrate the following:
Ablation: attention + flag. This model exhibited the worst performance, consistently underperforming throughout training compared to the other configurations. The absence of both the attention layer and flag bit limits the ability of the model to utilize shared information across tasks or explicitly distinguish between tasks. This indicates that both mechanisms are crucial for enabling effective multi-task learning.
Ablation: attention. Incorporating only the flag bit improved performance compared to the previously ablated model. The flag bit enables task differentiation by augmenting task representations with explicit task-specific information. However, the lack of a shared attention mechanism reduces the model’s ability to capture dynamic task relationships, resulting in suboptimal performance compared to models with attention.
Ablation: flag. This configuration, which retains the shared attention layer but removes the flag bit, outperformed the previous two models. The attention mechanism facilitates dynamic interaction across tasks, allowing the model to better leverage shared information, even in the absence of explicit task differentiation through flag bits.
Ablation: none (ours). The original configuration, which combines both the shared attention layer and the flag bit, achieved the best performance across all training epochs. This result highlights the complementary nature of the two components: the attention layer dynamically learns shared task interactions, while the flag bit provides explicit task differentiation. Together, these mechanisms enable the model to outperform all other configurations consistently, starting from the early stages of training.
VI Conclusion

### v2

@@ 5.5 LP-based Interleaved Learning (IMTL-LP) Results
Figure 4: Task-specific model performance is shown. It can be seen that LP-based method surpasses the other baselines on both push and stack tasks, with a higher significance in the latter. However, for the hit task, the improvement was constrained to the learning period around 400$\approx$ 400≈ 400100010001000.
Task-wise Analysis. To better understand the learning dynamics, we examine the performance of individual tasks throughout the learning process. As illustrated in Figure 4, the performance on both the Push and Stack tasks improves when all tasks are learned simultaneously, with a marked enhancement observed when using the proposed model. Notably, the Stack task exhibits a significant performance gap between our proposed model, IMTL-LP, and the other two baselines. For the Hit task, as in the other task we see an early reduction in prediction error with IMTL-LP but IMTL-RAND also starts to perform competitive in the second half of training. Table 2 shows the total iterations allocated for each task due to autonomous task selection. A higher performance can be obtained by IMTL-LP compared to IMTL-RAND in stack task even though it is allocated a shorter training time in total, which indicates that the amount of total training time allocated is not a direct determinant of task learning performance.
Model Push Hit Stack
IMTL-RAND 1002.5 ±24.4$\pm$ 24.4± 24.4 997.7 ±19.9$\pm$ 19.9± 19.9 1008.8 ±16.8$\pm$ 16.8± 16.8
IMTL-LP 1525.6 ±124.0$\pm$ 124.0± 124.0 624.8 ±93.2$\pm$ 93.2± 93.2 850.6 ±185.6$\pm$ 185.6± 185.6
Table 2: The average total number of training iterations received by each task as the result of autonomous task selection in IMTL-LP is compared with uniformly random task selection (IMTL-RAND). The average and standard deviation is over 10 training sessions.
5.6 Effect of Network Complexity on Learning Performance
Figure 5: Learning performance of baselines (SINGLE, IMTL-RAND) and our model (IMTL-LP) in different network complexity configurations. Results are taken at mid-training.
We analyzed the impact of network complexity on model performance by evaluating SINGLE, IMTL-RAND, and our proposed IMTL-LP (no energy modulation) at three levels of complexity: low (800 parameters), medium (2000 parameters), and high (5200 parameters), as shown in Figure 5. At low complexity, IMTL-LP significantly outperformed both baselines, highlighting its efficiency in resource-constrained conditions due to our learning progress-based interleaving strategy. At medium complexity, IMTL-LP still outperformed the baselines, although the gap narrowed, suggesting that while SINGLE and IMTL-RAND benefited from additional parameters, IMTL-LP maintained its advantage through better task management. At high complexity, the differences among the three models further decreased, with similar performance across models, indicating that increased capacity partially compensates for limitations in baseline strategies. However, IMTL-LP consistently remained competitive or best performing across all complexity levels.
5.7 Interleaved vs. Blocked Learning
Figure 6: Interleaved learning versus blocked learning prediction error plots for each task order combination. Each figure has three regions divided by vertical dashed lines, indicating where the BLOCK model switches to the next task in the order. P stands for Push, H stands for Hit and S stands for Stack task. Rightmost region in each figure shows where both models, BLOCK and IMTL-LP, have seen all the tasks to be learned. In all the task orders, IMTL-LP surpasses block learning.
Blocked learning is a training paradigm in which tasks are presented sequentially in distinct, uninterrupted intervals. It can be questioned that the blocked learning scheme may match or surpass the performance observed with interleaved learning when a certain task order is used. To test for this, we conducted a detailed comparison of interleaved learning and blocked learning with a neural network suitable for multi-task learning across the tasks of Push, Hit, and Stack introduced in Section 4.
Given the three tasks, there are six possible training orders (e.g., Push $\to$ Hit $\to$ Stack or Push $\to$ Stack $\to$ Hit, etc.). To see whether any order can perform similar or bettered that interleaved learning, we trained the BLOCK model across all six task orders (task IDs are denoted in the order as -XXX, e.g., BLOCK-123), while the IMTL-LP model remained unchanged across these configurations, as it does not depend on the task order (see Table 2 for per task training count of IMTL-LP model).
@@ 5.6 Effect of Network Complexity on Learning Performance
Figure 6 shows the performance of the IMTL-LP model and the BLOCK model for each of the six task orders. It can be seen that the proposed model, IMTL-LP consistently outperforms the BLOCK model across all task orders. No matter the order in which the tasks are trained, our IMTL-LP model always surpasses the blocked learning method, which shows that interleaved learning prevents catastrophic forgetting of learned information by switching the context in periods, thus being able to recall faster. On the other hand, blocked learning as expected suffers from catastrophic forgetting as evidenced by the error spikes at the task switch boundaries. Additionally, the variability in performance across different task orders for the BLOCK model indicates its sensitivity to the order of training.
@@
Figure 7: Comparison of baselines and LP-based model (IMTL-LP) to the EMLP-based model (IMTL-EMLP) where the sensitivity constant is set to $k=1$.
5.8 Modulation of LP with Computational Energy (IMTL-EMLP)
In this section, inspired form human brain’s tendency to reduce computational cost, we explore discounting the learning progress of the tasks by the neural cost they incur during learning. For this, we first test IMTL-EMLP that uses task selection according to $e^{k\cdot\tilde{LP}^{\mathcal{T}_{i}}}/\tilde{EC}^{\mathcal{T}_{i}}$ where the relative importance of LP set as $k=1$ (see Equation 9 ) and compare it to IMTL-LP and other baselines. Then we run experiments with different values of $k$ and compare the resulting learning performances in terms of prediction error and energy expenditure. Figure 7 presents the overall learning performance of our proposed interleaved learning (IMTL-LP and IMTL-EMLP) with the baseline models (SINGLE and IMTL-RAND), in terms of prediction performance. We can clearly say that our energy-modulated model, IMTL-EMLP, shows the best performance until 1000 epochs, being slightly better than our model with no energy modulation (IMTL-LP). After that, both of them show similar performance. The IMTL-EMLP model also outperforms both SINGLE and IMTL-RAND baselines, indicating that incorporating energy awareness does not nullify the gains of our architecture. The SINGLE model exhibits the weakest performance overall, reaffirming that task isolation limits skill learning efficiency. Meanwhile, IMTL-RAND shows modest improvements over SINGLE, highlighting that even random task interleaving can offer some benefit, yet remains clearly inferior to structured LP-based scheduling. Taken together, these results confirm that both learning progress and energy-aware arbitration strategies provide meaningful gains in training efficiency and final task performance, with IMTL-LP and IMTL-EMLP variants, converging faster and reaching lower prediction errors than all baselines.
@@ 5.8 Modulation of LP with Computational Energy (IMTL-EMLP)
To further evaluate the effect of EMLP-based task scheduling, we investigate the performance of the IMTL-EMLP model under varying values of the energy sensitivity coefficient $k$, and compare it to IMTL-LP and SINGLE baselines. As shown in Figure 8(a), which reports the MAE at the midpoint of the training (1500$\sim$ 1500∼ 1500), IMTL-LP achieves the lowest prediction error overall, confirming the strong effectiveness of LP-based task arbitration. Then, IMTL-EMLP-K=1.2 closely follows, demonstrating that modest energy modulation can preserve much of the learning performance. As $k$ decreases, the prediction performance gradually degrades, indicating that overly prioritizing energy conservation (i.e., smaller $k$) can hinder learning effectiveness. On the other hand, Figure 8(b) shows total energy consumption measured in terms of cumulative neuron activations. Here, we observe the opposite trend: energy usage decreases steadily with smaller $k$, with IMTL-EMLP-K=0.4 consuming the least energy overall, almost the same as the SINGLE baseline. While IMTL-LP achieves the best predictive performance, it is also the most energy-intensive. Notably, IMTL-EMLP-K=1.2 and K=1 provide a favorable trade-off, offering reductions in energy consumption with only a small sacrifice in prediction accuracy. This tunable behavior highlights the flexibility of the IMTL-EMLP framework, enabling agents to balance performance and energy efficiency according to task demands and environmental constraints.
@@
Figure 8: Effect of varying the energy sensitivity coefficient $k$ in the IMTL-EMLP model on (a) prediction performance and (b) total energy consumption, at the midpoint of training. Higher values of $k$ prioritize learning progress over energy efficiency, leading to lower prediction error but increased energy usage, while lower $k$ values reduce energy consumption at the cost of predictive accuracy.
5.9 Ablation Experiments
In order to reveal the individual affects of the self-attention and the task flag introduced in the proposed multitask learning architecture, a series of ablation experiments are conducted. In addition, task skill transfer ablations are conducted to investigate the skill transfer relations formed due to interleaved learning.
5.9.1 Self-Attention and the Task Flag Ablations
@@
Figure 9: Contributions of the shared attention and flag bit from the proposed multi-task learning architecture to the model performances. Neither attention nor flag alone can surpass the convergence speed of our proposed model, which combines them.
@@
The individual contributions of the attention layer and the flag bit in our proposed model are evaluated using four versions of the model: full model, without task flag, without attention layer, without flag and attention layer. Each model was trained using 10 random seeds. Below, we describe the four ablation settings and provide a detailed analysis of their performance:
Full model. The proposed model including a shared attention layer that process the task representations with the flag bit.
@@
Model w/o task flag. This variation retains the shared attention layer but removes the flag bits. Task representations are combined and passed to the shared attention layer without explicit differentiation using flags.
@@
Model w/o attention layer. In this model, the flag bits are concatenated to the task representations, same as the original matrix $Z$ from Equation (6), but instead of the shared attention layer, $Z$ is forwarded directly to task decoder of the current training task.
@@
Model w/o attention layer and flag. This configuration combines the two ablations thus only concatenated task representations are relayed to the task decoder of the current training task.
The results of the ablation experiments are shown together in Figure 9 as prediction error versus training epoch curves. We can deduce the following from the ablation experiments. The model without both the attention mechanism and the task-identifying flag bit (Figure 9, w/o attn&flag) showed the poorest performance, indicating that the absence of both components severely hinders the model’s ability to share information across tasks and distinguish between them. Introducing only the flag bit (Figure 9, w/o attn) improved the results, as it allowed for explicit task differentiation, though the lack of attention limited the model’s capacity to capture dynamic task relationships. Conversely, using the attention mechanism without the flag (Figure 9, w/o flag ) led to better performance, as the model could exploit shared information across tasks despite the absence of explicit task cues. The full model, incorporating both the attention layer and the flag bit, consistently outperformed all other configurations, demonstrating that the two mechanisms are complementary and together provide significant benefits for multi-task learning from the early stages of training.
5.9.2 Object-wise Skill Transfer Analysis
Figure 10: Impact of source-to-target task ablations on target task performance across objects. Each sub-figure corresponds to a specific object and visualizes the performance degradation (increase in loss) for target tasks (columns) when the representation of a particular source task (rows) is removed from the key and query inputs of the shared multi-head attention module. This experiment reveals the dependency of each target task on the shared representation of other source tasks, highlighting the inter-task transfer and interference dynamics in the multi-effect prediction framework.
In this section, we present skill transfer among tasks that were quantified by ablating task-to-task projections in the shared attention layer after training is completed. To assess a source task’s contribution to a target task, we define the loss obtained on the target task when all task representations are intact in the key/value matrices as $\mathcal{L}_{\mathrm{full}}$ and use $\mathcal{L}_{\mathrm{ablate}}$ to denote the loss obtained after zeroing out the representation of the source task. If the source task is irrelevant for the target task, a large increase in the loss is not expected. On the other hand, if the source task is utilized, then the loss should increase. With this logic, we define $\Delta\mathcal{L}=\mathcal{L}_{\mathrm{ablate}}-\mathcal{L}_{\mathrm{full}}$ and interpret the results based on this metric. For example, given a (source, target) task pair, a positive $\Delta\mathcal{L}$ indicates positive transfer from the source task to the target task. For skill transfer analysis, we focused on our IMTL-LP model and repeated training & testing 10 times with random seeds. The average impact of the ablations are measured by $\Delta\mathcal{L}$ and its standard deviation and used for numeric and visual representation (see Figure 10).
The skill transfer analysis is carried out per-object basis; so, the reported average $\Delta\mathcal{L}$ for a source and target task pair indicates the skill transfer between the tasks when a particular object is being handled.
First, we present the results for single-object interactions (Push and Hit), and then look at two-object interactions (Stack). Figure 10(a) shows skill transfer levels between source tasks (columns) and target tasks (rows) as a color coded matrix. During interaction with sphere, skill transfer was lower than interacting with the other objects. With sphere, the highest $\Delta\mathcal{L}$ is observed in $Stack\rightarrow Hit$ case with 0.860.860.86, whereas the lowest is observed in $Stack\rightarrow Push$ (0.220.220.22). Hit seems to better exploit Stack compared to Push suggesting that richer dynamics of multiple reflections from the walls (due to hit action) and the falls during unstable stacking shares shares certain computational similarities. Supporting this interpretation, in the horizontal cylinder a similar pattern can be observed where Stack benefits Hit more than Push. The common rollability affordance of sphere (2D) and the horizontal cylinder (1D) seem to limit the skill transfer among tasks when the interacted object is rollable. This may be explained by the fact that the behavior variety one can observe from rollable objects is higher compared to non-rollable objects, thereby making consistent skill transfer difficult. One difference in the sphere and the horizontal cylinder is the relative high skill transfer observed in $Push\rightarrow Hit$. This can be understood by considering that the push action is applied from arbitrary directions with many of them not inducing a rolling behavior. In those cases, the behavior of the object is closer to a stable object such as cube or horizontal-prism and thus the transfer level is somewhat shifted toward higher levels as in the stable objects (see horizontal prism and cube $Push\rightarrow Hit$ skill transfers). The effects of affordance and orientation can be seen in the skill transfer level of the cylinder and the prism: the $Push\rightarrow Hit$ transfer is limited when the objects are vertical (cylinder: 0.550.550.55; prism: 0.280.280.28), whereas the transfer is enhanced when they are placed horizontally (cylinder: 1.551.551.55; prism: 3.263.263.26). It seems that Hit action becomes beneficial to Push when objects are oriented vertical as opposed to horizontal so that they are knock-able. In the cylinder case, altering the orientation from horizontal to vertical causes a transfer change of 1.151.151.15 to 1.851.851.85; similarly in the prism case, a change of 3.073.073.07 to 3.333.333.33 is observed.
Moving on to the two-object interactions in Figure 10(b) for the Stack task, each matrix is titled with the base object on to which the picked object is placed. Rows show the picked object and columns show the source tasks (Push and Hit). The most prominent observation is that Hit task is strongly utilized by the Stack task compared to the Push task that have a non-negative small contribution. The high transfers level ($\Delta\mathcal{L}\approx 3.7-5.2$) from Hit task are observed when the object the base object stays intact, as in when the pair can be stacked successfully. The transfer levels when the base object is a sphere or a horizontal cylinder follow a similar pattern to that of one-object interactions. Their rollability affordances in 2D and 1D, respectively, cause them to behave differently depending on whether they receive a low impact force (push) or high impact force (hit). Cube or vertical cylinder as the base object show almost the same skill transfer pattern.
Considering these object-wise analyzes, along with the task-specific learning performance figures (see Figure 4), we see that Push task learns the effect prediction problem mostly on its own during the early stages of learning, while Hit and Stack tasks co-learn and benefit from each other throughout the training. It is intriguing to see that even though the task selection and learning performance is affected by the random initializations, the standard deviation of the effects of the ablations is about 14%. This indicates that although different skill transfer patterns are possible, the current results reflect a general pattern for the tasks addressed. Overall, we can conclude that, our task agnostic task selection mechanism, albeit being not geared to elicit skill transfer, sculpts an emergent transfer landscape that leads to faster and improved learning.
6 Conclusion